A certain automobile average 18 kilometers per liter on the highway. How many liters of gasoline are required to complete a 450 kilometer trip?
Answer by stanbon(75887)

A certain automobile average 18 kilometers per liter on the highway. How many liters of gasoline are required to complete a 450 kilometer trip?
---------------
Ans:: (450 km)/(18 km/liter) = (450/18) liters = 150/6 = 75/3 = 25 liters
-----
Cheers,
Stan H.
--------------